# NdLinear: Don’t Flatten! Building Superior Neural Architectures by Preserving N-D Structure

### v2

@@ 5 Conclusion
This paper introduced NdLinear, a novel linear transformation layer meticulously designed to address the prevalent issue of destructive input flattening. By operating directly on N-D tensors and applying transformations sequentially along each dimension, NdLinear fundamentally preserves the rich, inherent structural integrity of multi-dimensional data. This offers a robust alternative to conventional linear layers that often discard critical cross-axis information, thereby enabling a more faithful representation of N-D data throughout neural network processing.
Our comprehensive experiments across diverse domains as computer vision, natural language processing, time series, and tabular data, consistently demonstrated NdLinear’s compelling advantages. It achieves substantial parameter efficiency, with reductions ranging from valuable percentages in transformer blocks to over 90% in contexts like CNN feature extractors or compact classification heads. This efficiency is achieved while typically maintaining or, crucially, enhancing model performance. These improvements, observed across a variety of architectures (CNNs, RNNs, Transformers, MLPs), underscore NdLinear’s generality as a drop-in replacement. These empirical results strongly affirm our analysis regarding NdLinear’s beneficial inductive bias and efficiency gains.
By embedding N-D structure awareness into a foundational layer, NdLinear encourages a paradigm shift in designing neural architectures, moving from attempts to recover lost data relationships towards actively preserving them. This principled approach makes NdLinear a valuable component for building more capable and efficient large-scale AI systems, potentially reducing the need for certain compensatory preprocessing steps. The demonstrated strengths in efficiency and structure-aware representation learning open promising avenues in computationally intensive domains such as generative AI, complex scientific discovery (e.g., in genomics and multimodal data analysis), and in developing richer representations for intelligent agents, all of which warrant further exploration.
In summary, NdLinear provides a simple yet potent modification to a ubiquitous component of neural networks. By championing the preservation of N-D data structures as its core principle—to process data in its native form rather than flatten it—NdLinear offers a foundational advance towards more efficient, structurally-aware, and ultimately more effective AI models capable of better navigating our multi-dimensional world.
Appendix A Comparison with Structured Representation Learning Methods
We now compare NdLinear to other approaches that aim to capture structured dependencies or reduce parameters in high-dimensional models:
Graph Neural Networks (GNNs).
GNNs handle graph-structured data [Scarselli et al., 2008, Micheli, 2009, Bronstein et al., 2017, Zhou et al., 2020, Wu et al., 2020]. One can represent an $n$-dimensional grid (e.g., an image) as a graph, with each element as a node connected to its neighbors. GNNs often use graph convolutions or message passing [Gilmer et al., 2017, Kipf and Welling, 2017, Hamilton et al., 2017], capturing local structure without flattening. They aggregate features from neighboring nodes, acting like a convolution or local smoothing.
However, using GNNs for grid data can be overkill. They add significant computational overhead and do not leverage the regular structure. To reach global interactions, GNNs usually need many layers or long-range edges [Battaglia et al., 2018]. By contrast, NdLinear applies a global linear interaction along each dimension in one layer. A GNN typically shares weights across edges or edge types, so parameters do not necessarily grow with input size. Still, it may require many layers to propagate information globally. NdLinear instead uses the grid’s structure to apply dimension-wise interactions in one step, making it simpler and more efficient for tensor grids. GNNs excel at arbitrary graphs where no regular structure exists. Another difference is weight sharing: NdLinear shares weights fully along each dimension, while GNNs can have more complex or no sharing.
In summary, NdLinear is specialized for tensor grids and is faster and simpler on those tasks, whereas GNNs are more flexible but heavier. Hence, one typically does not replace a fully-connected layer for images with a GNN, due to the overhead.
Tensor Decomposition Methods.
NdLinear is closely related to tensor decomposition methods used for model compression. Examples include Tucker [Tucker, 1966], CP (CANDECOMP/PARAFAC) [Carroll and Chang, 1970, Harshman, 1970], and Tensor-Train (TT) Decompositions [Oseledets, 2011, Novikov et al., 2015]. These methods view a large weight matrix (like in fully-connected layers) as a high-order tensor. They factorize this tensor into smaller components. For instance, the Tensor-Train decomposition represents weights as sequences of smaller tensors. This achieves significant parameter reduction with minimal accuracy loss [Novikov et al., 2015].
@@ Appendix A Comparison with Structured Representation Learning Methods
NdLinear can be seen as a hand-crafted factorization of a fully-connected weight matrix. Specifically, the full weight matrix $W_{\text{full}}$ implicitly has a Kronecker product structure derived from mode-wise matrices $\{W_{1},\dots,W_{n}\}$. This corresponds to a rank-1 Tucker decomposition without a core tensor (or equivalently, a core of rank 1 in each mode).
@@
The main trade-off here is expressiveness vs. efficiency. NdLinear’s decomposition is low-rank in a multilinear sense. It cannot represent unique, non-factorizable interactions between dimensions. More flexible decompositions, such as full Tucker or higher-rank tensor decompositions, capture more interactions. However, these approaches require significantly more parameters compared to NdLinear’s simple sum $\sum_{i}D_{i}H_{i}$. They can also be more challenging to train, sometimes needing special initialization or multi-stage training.
@@
To combat these, NdLinear balances these trade-offs well. It is highly parameter-efficient, easy to train from scratch, and captures essential structured dependencies. If needed, one can extend NdLinear by increasing the factorization rank. For example, one could learn multiple $W_{i}$ matrices per mode and sum their effects, analogous to a rank-$R$ core. However, our experiments show that the simple version already performs very well. Compared to general tensor decomposition methods, NdLinear is also more interpretable. Each $W_{i}$ clearly indicates how dimension $i$ is transformed, rather than dispersing transformations across multiple factors.
Grouped and Factorized Convolutions.
Various convolutional factorizations reduce CNN computation, notably:
  • Grouped convolutions [Krizhevsky et al., 2012a, Xie et al., 2017] restrict filters to subsets of channels, reducing parameters.
  • Depthwise separable convolutions [Howard et al., 2017a, Chollet, 2017a] split convolution into per-channel depthwise and pointwise ($1\times 1$) operations, significantly cutting FLOPs.
NdLinear applies a similar concept to fully-connected layers, decomposing weights by dimension rather than channel. Instead of one large weight matrix mapping $\prod D_{i}$ inputs to $\prod H_{i}$ outputs, NdLinear factorizes it into $n$ matrices, each mapping $D_{i}$ to $H_{i}$. This sequential dimension-wise grouping yields structured weights, improving efficiency with minimal accuracy loss.
Unlike grouped or separable convolutions that typically serve as internal layers, NdLinear naturally produces structured multi-dimensional outputs $(H_{1},\dots,H_{n})$, making it suitable for structured prediction tasks.
Other Specialized Layers.
Several specialized approaches for structured representation learning have been explored:
  • Slicing-based layers [Shao et al., 2016, Dieleman et al., 2016] separately process spatial, temporal, or rotated input segments, preserving structure without flattening.
  • Capsule Networks [Hinton et al., 2011, Sabour et al., 2017, Hinton et al., 2018] use vector or matrix capsules with routing mechanisms to maintain hierarchical relationships.
  • Hadamard/Fourier feature mixing [Rahimi and Recht, 2007, Le et al., 2013, Tancik et al., 2020, Lee-Thorp et al., 2022, Pan et al., 2022] employs fixed, non-learnable transforms (e.g., Walsh-Hadamard, Fourier) for global mixing.
NdLinear is simpler and more flexible than these methods, applying learned factorized linear transformations along each dimension without specialized routing or static transforms.
@@ Appendix B More Related Work
Deep neural networks often contain millions of parameters, many of which are highly redundant. For instance, Denil et al. [2013] demonstrate that a large fraction of weights can be predicted from a small subset. In some cases, 95% of parameters need not be learned with no loss in accuracy. This redundancy has motivated numerous efficient parameterization strategies that preserve model expressiveness while drastically reducing storage and computation requirements.
@@
A prominent research direction exploits high-order (multi-dimensional) structures of weight tensors using tensor factorization. Lebedev et al. [2015] apply CP decomposition to 4D convolutional kernels, reducing parameters and inference cost. Novikov et al. [2015] introduce Tensor Train (TT) layers, compressing fully-connected layers into compact tensors without losing expressiveness. More recently, Ye et al. [2020] propose Block-Term Tensor Networks, approximating network weights through block-term (Tucker) decompositions. These tensor-structured layers significantly reduce parameters and preserve high-order interactions.
@@
Other methods impose algebraic structure on weight matrices or use parameter sharing to reduce model size. Sindhwani et al. [2015] replace dense matrices with Toeplitz-like structures, significantly compressing layers without losing performance. Similar methods include circulant matrices, block-circulant transforms, and low-rank factorizations [Lebedev et al., 2015]. These structured layers consistently match or exceed unstructured models in performance, using fewer parameters and lower computational cost.
@@
Emerging multi-space learning methods embed data into multiple latent spaces to preserve complex relationships. Rather than relying only on Euclidean spaces, these methods project data into multiple spaces (e.g., Euclidean and hyperbolic) to capture richer structures. Wang et al. [2023] embed LiDAR features in Euclidean and hyperbolic spaces, improving hierarchical information encoding and pose regression. Multi-space learning complements tensor factorization by enhancing model expressiveness without increasing layer size.
@@
A common theme across these efforts is that preserving high-order structure leads to compact yet expressive representations. Viewing weight or activation tensors in their natural multi-dimensional format directly captures interactions along each dimension. Depthwise separable convolutions [Chollet, 2017b, Howard et al., 2017b] partially achieve this by separating channel-wise and spatial-wise operations. However, most fully-connected layers still flatten inputs, ignoring multi-axis dependencies.
The proposed NdLinear approach naturally fits within structured methods. NdLinear iterates through each dimension of an $n$-dimensional tensor, applying dimension-specific linear transforms. It preserves multi-axis relationships without extra computational cost. Similar to depthwise separable convolutions, NdLinear generalizes efficiently to more dimensions and tasks beyond convolutional features. Combining NdLinear with existing factorization methods can further enhance multi-space learning in large-scale models.
These advances in structured tensor methods, multi-space feature learning, and efficient parameterization show that deep networks can be both lightweight and expressive. Exploiting algebraic structures or multiple embedding spaces yields compact models, often with improved accuracy—essential for large-scale or resource-constrained applications.
Appendix C Implementation Details and Training Protocol

### v3

@@ 5 Conclusion
We introduced NdLinear, a linear layer that operates directly on N-dimensional tensors through sequential dimension-wise transformations. This simple change, by processing tensors in their native form rather than flattening them, yields dramatic improvements across modern deep learning.
Our extensive experiments demonstrate that NdLinear reduces parameters by 50-95% while maintaining or improving performance across all tested configurations. The results reveal striking consistency: from NdLinear-LoRA fine-tuning LLaMA-8B with 9× fewer parameters while improving reasoning accuracy, to achieving 70% lower error in time-series forecasting with half the parameters, to outperforming alternative structured methods (TRL/TCL, TT). This universality, spanning language, vision, time-series, and tabular domains, validates dimension-wise factorization as a fundamental principle.
Theoretically, we proved NdLinear maintains expressivity through preserved VC-dimension scaling, explaining why dramatic parameter reduction doesn’t sacrifice performance. Our controlled experiments precisely quantified when the method’s inductive bias helps versus hurts: NdLinear excels when data exhibits axis-separable structure, but struggles with highly entangled patterns. This theoretical understanding, combined with our ablations showing robustness to hyperparameters and negligible computational overhead (<2%), provides clear deployment guidance.
Beyond its immediate practical benefits, NdLinear challenges a fundamental assumption in neural architecture design: that flattening is necessary for linear transformations. By demonstrating that structure-preserving operations consistently outperform their flattened counterparts, we reveal that the ubiquitous practice of tensor flattening has been systematically discarding valuable inductive biases. As neural networks evolve to process increasingly complex multidimensional data, from volumetric medical imaging to spatiotemporal climate models, NdLinear’s principle of native dimensional processing offers not just an optimization, but a paradigm shift toward architectures that inherently respect and leverage the structure of our multidimensional world.
Appendix A Detailed Comparison with Related Methods
NdLinear’s design as a structure-preserving, parameter-efficient linear layer for N-D tensors is informed by, yet distinct from, several established concepts in machine learning and tensor algebra. We detail these relationships below.
A.1 Tensor Decomposition Methods
NdLinear’s core mechanism employs mode-wise tensor-matrix products prominent in Tucker decomposition [Tucker, 1966, Kolda and Bader, 2009]. However, unlike classical tensor decomposition methods that primarily analyze or compress static data or pre-existing weight tensors [Novikov et al., 2015, Newman et al., 2024], NdLinear integrates these mode-wise operations as a learnable, dynamic layer within a network. Its purpose is efficient forward transformation of activations while preserving N-D structure, not data analysis or post-hoc model compression.
@@ A.1 Tensor Decomposition Methods
NdLinear can be viewed as a hand-crafted factorization of a fully-connected weight matrix. Specifically, the full weight matrix $W_{\text{full}}$ implicitly has a Kronecker product structure derived from mode-wise matrices $\{W_{1},\ldots,W_{n}\}$. This corresponds to a rank-1 Tucker decomposition without a core tensor (or equivalently, a core of rank 1 in each mode).
@@
The main trade-off is expressiveness vs. efficiency. NdLinear’s decomposition is low-rank in a multilinear sense—it cannot represent arbitrary non-factorizable interactions between dimensions. More flexible decompositions (full Tucker or higher-rank tensor decompositions) capture more interactions but require significantly more parameters compared to NdLinear’s simple sum $\sum_{i}D_{i}H_{i}$ and can be more challenging to train, sometimes needing special initialization or multi-stage training.
@@
If needed, one can extend NdLinear by increasing the factorization rank (e.g., learning multiple $W_{i}$ matrices per mode and summing their effects, analogous to a rank-$R$ core). However, our experiments show the simple rank-1 version already performs well. Each $W_{i}$ clearly indicates how dimension $i$ is transformed, providing better interpretability than general tensor decomposition methods that disperse transformations across multiple factors.
A.2 Factorized Convolutions and Axial Operations
Neural nets have long exploited axis-wise structure to cut parameters and impose useful priors. In CNNs, an image tensor $X\!\in\!\mathbb{R}^{H\times W\times C}$ is not flattened; instead, convolutions slide local kernels over the two spatial axes $(H,W)$ while mixing channels $C$. In sequence models, a tensor $X\!\in\!\mathbb{R}^{L\times C}$ (length $L$, features $C$) is processed along the sequence axis $L$ (e.g., attention/conv) while mixing feature channels $C$ within positions. Below we recap common factorized operators and their limitations.
Definitions.
Channels are feature maps carried alongside a position-like axis (e.g., RGB planes or intermediate filters in CNNs; hidden features per token in sequences). Sequence (or spatial) dimensions are position-like axes along which locality or ordering matters (e.g., time $L$, image height $H$, width $W$).
Grouped convolutions.
Grouped convs restrict each filter to operate on a subset of channels, partitioning $C$ into groups to reduce parameters and FLOPs [Krizhevsky et al., 2012, Xie et al., 2017]. Limitation: channel mixing is constrained within groups; cross-group interactions require additional layers, and the grouping choice is a manual architectural prior.
Depthwise separable convolutions.
Depthwise separable convs factor a full conv into a depthwise spatial conv per channel followed by a $1{\times}1$ (pointwise) conv that mixes channels [Chollet, 2017b, Howard et al., 2017]. This yields large FLOP/parameter savings with strong accuracy. Limitations: the factorization is tied to 2D spatial structure; expressivity hinges on the pointwise mixer; extending beyond standard spatial axes typically needs custom kernels.
Axial (factorized) attention.
Axial attention decomposes 2D/3D attention into a sequence of 1D attentions applied along one axis at a time (e.g., height then width), cutting quadratic costs while preserving long-range interactions along each axis [Ho et al., 2019, Wang et al., 2020, Yan et al., 2023]. Limitations: axis order becomes an architectural prior; full joint interactions across axes emerge only after stacking, and costs can still be high for very long axes.
CNNs.
CNNs are the canonical instance of axis-aware processing: they exploit spatial locality (factorization in $(H,W)$) and defer heavy channel mixing to $1{\times}1$ pointwise layers. Grouped/depthwise variants intensify this factorization to further reduce compute.
NdLinear in this landscape.
NdLinear generalizes the factorization principle beyond conv/attention mechanics: given an $n$-D activation $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{n}}$, it systematically factorizes a linear map into mode-wise transforms $D_{i}\!\to\!H_{i}$ and applies them sequentially, producing a structured output $Y\!\in\!\mathbb{R}^{H_{1}\times\cdots\times H_{n}}$. Thus, instead of a single dense matrix $\mathbb{R}^{\prod_{i}D_{i}}\!\to\!\mathbb{R}^{\prod_{i}H_{i}}$, NdLinear uses $n$ small matrices with $\sum_{i}D_{i}H_{i}$ parameters, preserving the N-D shape throughout. Unlike grouped/depthwise convs (tied to spatial kernels) or axial attention (tied to attention mechanics and axis ordering), NdLinear: (i) applies to arbitrary N-D tensors (images, videos, spectrograms, multivariate time series, tabular tensors), (ii) naturally supports both compression and expansion per mode $(H_{i}\lessgtr D_{i})$, (iii) creates insertion points for normalization/activation between mode maps, and (iv) preserves a clear axis-wise inductive bias without requiring handcrafted groups or convolutional kernels.
Known failure modes and how NdLinear relates.
Axis-factorized operators can underperform when tasks require strong entangled cross-axis interactions (e.g., patterns tied jointly to $(H,W)$ rather than separably to each). Grouped/depthwise convs may also struggle if the chosen grouping misaligns with semantics; axial attention can be sensitive to axis order and depth. NdLinear shares the core trade-off (axis-wise separability vs. full expressivity) but makes it explicit and tunable by per-mode widths/ranks or stacking; it inherits the efficiency benefits of factorization while remaining N-D agnostic in form.
A.3 Tensor Contraction and Regression Layers
Tensor Contraction Layers (TCL).
TCLs [Cichocki, 2014, Novikov et al., 2015] implement a one-shot multilinear map by contracting an input tensor $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{m}}$ with a set of mode matrices $\{V^{(k)}\in\mathbb{R}^{R_{k}\times D_{k}}\}_{k=1}^{m}$, yielding a reduced tensor $X\times_{1}V^{(1)}\times_{2}\cdots\times_{m}V^{(m)}\in\mathbb{R}^{R_{1}\times\cdots\times R_{m}}$. The goal is typically dimensionality reduction (feature compression) via tensor contractions prior to downstream layers; intermediate positions for normalization/activation and per-mode biases are usually not part of the basic formulation [Cichocki, 2014, Novikov et al., 2015].
Tensor Regression Layers (TRL).
TRLs [Kossaifi et al., 2020] cast prediction as supervised regression/classification with a fixed tensor format for the weight (e.g., Tucker-/TT-structured). Given an input tensor $X$, a TRL fits a low-rank tensor $W$ (plus optional bias) such that $\langle W,X\rangle$ (or a nonlinear variant) matches targets. The emphasis is on learning with low-rank weights for sample/parameter efficiency; the tensor format (ranks/cores) is chosen a priori and does not expose interleaved normalization/activation between modes [Kossaifi et al., 2020].
NdLinear vs. TCL/TRL.
NdLinear is a structure-preserving, learnable linear layer that applies sequential mode-wise maps to activations, $Y\;=\;X\times_{1}W_{1}\times_{2}\cdots\times_{m}W_{m}\in\mathbb{R}^{H_{1}\times\cdots\times H_{m}},$ with an implicit rank-1 Tucker/Kronecker weight and $\sum_{k}D_{k}H_{k}$ parameters. This design differs in intent and mechanics:
  • Goal. TCL primarily contracts modes for reduction [Cichocki, 2014, Novikov et al., 2015]; TRL fixes a low-rank regression format [Kossaifi et al., 2020]; NdLinear is a drop-in linear alternative for N-D activations that preserves the full tensor shape and supports expansion or compression per mode $(H_{k}\gtrless D_{k})$.
  • Execution. TCL performs a single contraction; NdLinear performs sequential per-mode maps, creating natural insertion points for LayerNorm/Dropout/activations and allowing per-mode biases. TRL optimizes a fixed low-rank weight but does not expose interleaved, mode-by-mode transforms during the forward pass.
  • Expressivity/control. All three impose structured priors; NdLinear’s axis-wise separability can be tuned via widths/ranks or stacked blocks, interpolating between strong separability and near-dense behavior while retaining N-D outputs.
TCL as a special case of NdLinear.
Under the joint constraints (i) $H_{k}{=}R_{k}$ for all $k$ (no expansion beyond the contracted size), (ii) no per-mode biases, and (iii) no interleaved operations between mode maps (i.e., a single, commutative product), the one-shot TCL contraction coincides with NdLinear [Cichocki, 2014, Novikov et al., 2015]. Outside this narrow corner, NdLinear is strictly more general and practical: it preserves N-D structure through sequential maps, supports per-mode biases and interleaving (stability), and flexibly expands or compresses each mode.
A.4 Structured Linear Layers
NdLinear sits within the broader family of structured linear layers for parameter efficiency [Denil et al., 2013, Wei et al., 2024]. Classic approaches constrain a large dense weight $W\!\in\!\mathbb{R}^{(\prod_{i}D_{i})\times(\prod_{i}H_{i})}$ by imposing algebraic structure on $W$ itself—e.g., low-rank factorizations [Sainath et al., 2013], block/Butterfly/Monarch-style sparse–fast transforms [Dao et al., 2022], or other learned structured matrices [Sindhwani et al., 2015, Potapczynski et al., 2024]. These methods decouple the parameterization of $W$ from the native organization of the activations, often yielding strong compression but requiring the model to implicitly discover how that structure aligns with the data.
In contrast, NdLinear derives its structure from the N-D layout of the input. Given $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{n}}$, NdLinear replaces the dense map $\mathbb{R}^{\prod_{i}D_{i}}\!\to\!\mathbb{R}^{\prod_{i}H_{i}}$ with sequential mode-wise transforms $D_{i}\!\to\!H_{i}$, preserving tensor shape and inducing an explicit Kronecker (rank-1 Tucker) weight with only $\sum_{i}D_{i}H_{i}$ parameters. This data-centric factorization (i) makes the axis-wise inductive bias transparent and tunable (via per-mode widths/ranks or stacking), (ii) supports per-mode expansion or compression $(H_{i}\gtrless D_{i})$, and (iii) exposes natural insertion points for normalization/activation between modes—while retaining the efficiency benefits typical of structured linear layers [Denil et al., 2013, Wei et al., 2024, Sainath et al., 2013, Dao et al., 2022, Sindhwani et al., 2015, Potapczynski et al., 2024].
A.5 Graph Neural Networks
Graph Neural Networks (GNNs) address data with irregular connectivity by propagating information over edges via message passing [Scarselli et al., 2008, Micheli, 2009, Bronstein et al., 2017, Zhou et al., 2020, Wu et al., 2020]. Concretely, a GNN updates node features by aggregating messages from neighbors and applying learnable transforms [Gilmer et al., 2017, Kipf and Welling, 2017, Hamilton et al., 2017]. Stacking layers increases the receptive field and enables global interaction, but typically requires multiple rounds of propagation to mix distant nodes [Battaglia et al., 2018].
While an N-D grid (e.g., image or tensor lattice) can be modeled as a graph (one node per cell, edges to local neighbors), this introduces unnecessary overhead on regular grids: message passing is inherently local, so achieving global mixing along each axis often demands many layers, with added memory/compute and potential optimization issues (e.g., depth-related bottlenecks). Moreover, parameter sharing in GNNs is tied to edge types and neighborhood schemas, not directly to axis-wise tensor structure.
NdLinear takes the complementary route for regular tensor grids. Given $X\!\in\!\mathbb{R}^{D_{1}\times\cdots\times D_{n}}$, it applies global, mode-wise linear maps $D_{i}\!\to\!H_{i}$ in a single layer, mixing information along entire axes without constructing a graph or iterating local messages. This preserves the N-D layout, yields $\sum_{i}D_{i}H_{i}$ parameters via an implicit Kronecker (rank-1 Tucker) structure, and exposes insertion points for normalization/activation between mode maps. In short: GNNs excel when connectivity is irregular or non-Euclidean [Bronstein et al., 2017, Zhou et al., 2020], whereas NdLinear specializes in axis-aware, structure-preserving transformations on regular tensors, providing simpler and often faster global mixing along each dimension.
When to use which.
Use GNNs for arbitrary graphs, heterogeneous edge types, and relational reasoning on non-grid data [Scarselli et al., 2008, Micheli, 2009, Bronstein et al., 2017, Zhou et al., 2020, Wu et al., 2020]. Use NdLinear when data are naturally arranged as regular tensors and you want efficient, axis-wise global interactions without flattening; its bias toward axis separability provides parameter savings and predictable behavior on grid-structured domains.
A.6 Other Specialized Approaches
A variety of specialized architectures preserve structure without resorting to full flattening:
Slicing-based layers.
Methods that slice inputs along spatial/temporal (or rotated) subdomains process each slice with shared weights, then recombine [Shao et al., 2016, Dieleman et al., 2016]. This preserves locality and orientation information with modest compute. Limitations: boundaries between slices can hinder cross-slice interaction; designs are task-/geometry-specific and often require bespoke preprocessing.
Capsule Networks.
Capsules use vector/matrix-valued units and routing to model part–whole hierarchies and pose relationships [Hinton et al., 2011, Sabour et al., 2017, Hinton et al., 2018]. They maintain structured representations through learned agreement between capsules. Limitations: routing adds iterative, nontrivial overhead; scaling to large resolutions and datasets has proven challenging; design choices (routing, capsule size) are sensitive.
Hadamard/Fourier feature mixing.
Fixed orthogonal or Fourier-like transforms provide global mixing with $\mathcal{O}(N\log N)$ or even $\mathcal{O}(N)$ cost (e.g., Random Features, Fastfood, FNet, Block-based variants) [Rahimi and Recht, 2007, Le et al., 2013, Tancik et al., 2020, Lee-Thorp et al., 2022, Pan et al., 2022]. Limitations: transforms are fixed (non-learnable) or only weakly parameterized, so alignment with data structure must be recovered by subsequent layers; expressivity depends on depth.
Relation to NdLinear.
NdLinear applies learned, factorized linear maps along each tensor mode, preserving the full N-D layout while enabling axis-wise global mixing with $\sum_{i}D_{i}H_{i}$ parameters. Unlike slicing [Shao et al., 2016, Dieleman et al., 2016], it does not require hand-crafted partitions; unlike capsules [Hinton et al., 2011, Sabour et al., 2017, Hinton et al., 2018], it avoids iterative routing; unlike fixed Hadamard/Fourier mixers [Rahimi and Recht, 2007, Le et al., 2013, Tancik et al., 2020, Lee-Thorp et al., 2022, Pan et al., 2022], it learns mode-wise transforms end-to-end. This yields a simple, geometry-agnostic mechanism for structure-preserving, parameter-efficient linear transformation on regular tensors.
@@ Appendix B More Related Work
Modern neural networks contain substantial parameter redundancy: a large fraction of weights can be predicted from a small subset, sometimes up to 95% with no loss in accuracy [Denil et al., 2013]. This has motivated a broad line of work on efficient parameterizations that preserve accuracy while reducing storage and compute.
@@
A major thread leverages high-order structure via tensor decompositions of weights. CP/Tucker-style compressions applied to convolutional kernels reduce parameters and inference cost [Lebedev et al., 2015]. Tensor Train (TT) layers compress fully-connected mappings into compact tensorized operators [Novikov et al., 2015]. Block-Term (BT) tensor networks combine Tucker- and CP-like structure for additional flexibility [Ye et al., 2020]. These tensor-structured layers reduce parameters while retaining rich multi-way interactions by factoring weights across modes.
@@
Another approach imposes algebraic structure on large dense matrices, replacing them with families that admit fast transforms and fewer degrees of freedom. Toeplitz-like and related structured operators provide strong compression with competitive accuracy [Sindhwani et al., 2015]; related families (e.g., circulant, block-circulant) and low-rank factorizations likewise trade unrestricted expressivity for parameter/compute efficiency [Lebedev et al., 2015].
@@
Complementary to structural compression, multi-space learning embeds features into multiple geometries to better capture hierarchy and long-range relations. For example, jointly using Euclidean and hyperbolic spaces for LiDAR yields improved hierarchical encoding and pose estimation [Wang et al., 2023]. Such representations enhance expressivity without necessarily increasing individual layer sizes.
@@
Operational layers that respect native tensor axes often strike favorable accuracy–efficiency trade-offs. Depthwise separable convolutions split channel-wise and spatial mixing to cut FLOPs while preserving inductive bias [Chollet, 2017b, Howard et al., 2017]. However, many fully-connected stages still flatten activations, discarding axis structure learned upstream.
Positioning NdLinear.
NdLinear aligns with these trends but differs in where structure is imposed: rather than factorizing weights after flattening, it performs mode-wise learned linear maps directly on N-D activations, preserving tensor layout throughout. Conceptually, it is a rank-1 Tucker (Kronecker) parameterization of the dense linear map, with parameters scaling as $\sum_{i}D_{i}H_{i}$ rather than $\prod_{i}D_{i}\prod_{i}H_{i}$. This data-aligned factorization complements tensorized weights [Lebedev et al., 2015, Novikov et al., 2015, Ye et al., 2020] and structured matrices [Sindhwani et al., 2015], and, like depthwise separable convolutions [Chollet, 2017b, Howard et al., 2017], leverages axis-aware inductive bias—without resorting to flattening.
Appendix C Proofs and Technical Details
C.1 VC-Dimension Analysis
We analyze the expressive capacity of NdLinear compared to standard linear layers. Following Bartlett et al. [2019], any piecewise-linear feedforward network with $P$ parameters has VC-dimension $\Theta(P\log P)$.
Theorem C.1 (VC-Dimension of NdLinear).
Consider input tensors of shape $(B,a,b,c)$ transformed to outputs of shape $(B,d,d,d)$. Let:
  • $N_{\text{vanilla}}=dabc$ (parameters in vanilla linear layer)
  • $N_{\text{nd}}=d(a+b+c)$ (parameters in NdLinear)
Then:
  1. 1.
    As $d\to\infty$ with $a,b,c$ fixed: $N_{\text{nd}}=\Theta(N_{\text{vanilla}})$
  2. 2.
    NdLinear’s VC-dimension is $\Theta(N_{\text{nd}}\log N_{\text{nd}})$
Proof.
For part (1), observe that:
$N_{\text{nd}}=d(a+b+c)=\frac{a+b+c}{abc}\cdot dabc=\frac{a+b+c}{abc}\cdot N_{\text{vanilla}}.$
Since $\frac{a+b+c}{abc}$ is a positive constant (for fixed $a,b,c$), we have:
$N_{\text{nd}}=\Theta(N_{\text{vanilla}})\quad\text{as }d\to\infty.$
For part (2), by the Bartlett et al. result, since NdLinear has $N_{\text{nd}}$ parameters and maintains piecewise-linear structure through ReLU activations:
$\text{VCdim}_{\text{NdLinear}}=\Theta(N_{\text{nd}}\log N_{\text{nd}}).$
Interpretation: While NdLinear uses fewer parameters for finite $d$, as the hidden dimension grows, its parameter count becomes proportional to vanilla linear layers, preserving the same VC-dimension scaling.
Theorem C.2 (Parameter Count Lower Bound).
For positive integers $a,b,c,d$:
$d(a+b+c)>\log(dabc).$
Proof.
We have:
$\log(dabc)$ $=\log d+\log a+\log b+\log c$
$\leq(d-1)+(a-1)+(b-1)+(c-1)\quad\text{(using }\log x\leq x-1\text{)}$
$=d+a+b+c-4.$
Therefore:
$d(a+b+c)-\log(dabc)$ $\geq d(a+b+c)-(d+a+b+c-4)$
$=(d-1)(a+b+c)-d+1+4$
$=(d-1)(a+b+c-1)+4$
$\geq 4>0,$
since $d\geq 1$ and $a+b+c\geq 3$ for non-trivial tensors. ∎
C.2 Peak Activation Memory Analysis
Proposition C.1 (Peak Memory Overhead Bound).
For an input tensor with $m$ modes of sizes $(d_{1},\ldots,d_{m})$ where $\prod_{i=1}^{m}d_{i}=D$, and output sizes $k_{i}\leq d_{i}$, the peak additional activation memory for backpropagation satisfies:
$\frac{\text{extra activation memory}}{\text{baseline activation memory}}\leq\frac{\max_{i}(k_{i}/d_{i})\cdot\min_{i}d_{i}}{m\cdot\min_{i}d_{i}}\leq\frac{1}{m}.$
For typical 3D tensors (m=3), this overhead is at most 33%.
Proof.
During forward pass, NdLinear sequentially transforms each mode. For backpropagation, we store intermediate activations after each mode transformation.
After transforming $j$ modes, the tensor has shape:
$(B,k_{1},\ldots,k_{j},d_{j+1},\ldots,d_{m}).$
The peak extra memory occurs at the stage with the largest intermediate tensor. Since $k_{i}\leq d_{i}$, each intermediate tensor has at most $BD$ elements. The baseline memory is also $BD$ elements.
In the worst case where all $k_{i}=d_{i}$, we have at most (m1)(m-1) intermediate tensors to store, but only one is needed at any given time during backprop (due to sequential processing). Therefore:
$\text{Peak overhead}=\frac{BD\cdot\max_{i}(k_{i}/d_{i})}{BD}\leq\frac{1}{m}.$
C.3 Computational Complexity
Proposition C.2 (Exact FLOP Count).
NdLinear transforming $\mathcal{X}\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ to $\mathcal{Y}\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$ requires:
$\text{FLOPs}_{\text{NdLinear}}=2B\sum_{k=1}^{N}\left[\left(\prod_{j=1}^{k-1}H_{j}\right)\left(\prod_{j=k+1}^{N}D_{j}\right)D_{k}H_{k}\right],$
where the factor of 2 accounts for multiply-add operations.
This is typically $\mathcal{O}(BND^{N+1})$ for $D_{i}=H_{i}=D$, compared to $\mathcal{O}(BD^{2N})$ for vanilla linear layers—yielding orders of magnitude savings as $N$ increases.
Appendix D Implementation Details and Training Protocol